# How Far Can 100 Samples Go? Unlocking Overall Zero-Shot Multilingual Translation via Tiny Multi-Parallel Data

### 22 Jan 2024 v1

@@ Abstract
Zero-shot translation is an open problem, aiming to translate between language pairs unseen during training in Multilingual Machine Translation (MMT). A common, albeit resource-consuming, solution is to mine as many translation directions as possible to add to the parallel corpus. In this paper, we show that the zero-shot capability of an English-centric model can be easily enhanced by fine-tuning with a very small amount of multi-parallel data. For example, on the EC30 dataset, we show that up to +21.7 ChrF non-English overall improvements (870 directions) can be achieved by using only 100 multi-parallel samples, meanwhile preserving capability in English-centric directions. We further study the size effect of fine-tuning data and its transfer capabilities. Surprisingly, our empirical analysis shows that comparable overall improvements can be achieved even through fine-tuning in a small, randomly sampled direction set (10%). Also, the resulting non-English performance is quite close to the upper bound (complete translation). Due to its high efficiency and practicality, we encourage the community 1) to consider the use of the fine-tuning method as a strong baseline for zero-shot translation and 2) to construct more comprehensive and high-quality multi-parallel data to cover real-world demand.
@@ 1 Introduction
The zero-shot capability shown by Multilingual Machine Translation (MMT) Johnson et al. (2017) holds significant importance, particularly in the context of translations between low-resource or distant language pairs. However, even for systems trained on large-scale data, the zero-shot capability is still far from real-world demands (Tan and Monz, 2023), especially when scaling up the number of involved languages. To this end, substantial efforts (Zhang et al., 2020; Pan et al., 2021; Gu and Feng, 2022; Mao et al., 2023) have been dedicated to improving zero-shot capabilities of models trained on readily available, predominantly English-centric corpus.
@@
Figure 1: (a) English-centric training data is normally readily available but can only cover a few real-world directions, while (b) complete translation (Freitag and Firat, 2020) aims to cover all but suffers from the small data scale. (c) Mining partial non-English data as the bridge languages shows promising improvements but is also resource-consuming when scaling up. (d) We show that evident overall improvements can be achieved by fine-tuning an English-centric model with tiny extra multi-parallel data, which is readily available, like NTREX (Federmann et al., 2022).
@@
To fully cover translation directions, Freitag and Firat (2020) propose to mine multi-parallel (multi-way aligned) examples to extend the training set from English-centric to complete multilingual one as shown in Figure 1-(b). Non-English translation quality in this setting indeed increases a lot. However, such a setting is far away from real-world practice when scaling up: as shown in  (Freitag and Firat, 2020), to solely extend the training set from bilingual aligned to all languages (6-way) involved in their case, the amount of available data drops from 123M to 10K, which is insufficient.
@@
To reflect translation needs worldwide, Fan et al. (2021) build and open-source a training dataset covering 100 languages through industry-scale mining. In addition to English-centric data, supervised data for thousands of bridge language pairs is mined and included, organized based on language families. The MMT model trained on the resulting data, M2M100, exhibits evident non-English improvement in many non-English directions. This work drives a simple but also resource-consuming solution for real-world demand: mining as much training data as possible to bridge non-English language pairs, at least on the language family level.
@@
In this paper, we take a step back and again look at the readily available English-centric model. We empirically show that the corresponding zero-shot translation ability can be already unlocked via fine-tuning an English-centric model with a tiny amount of multi-parallel data, which is vastly simpler and more efficient than the extensive bridge data mining done by earlier work. Furthermore, we investigate the size effect of fine-tuning data: 1) Surprisingly, even if fine-tuning in randomly sampled 10% directions, the overall improvements are comparable to that of fully directional fine-tuning. 2) The improvements brought by very small fine-tuning datasets only slightly lag behind the upper bound (complete translation) while preserving English-centric capabilities, showing great practical potential. 3) Intriguingly, even with just one single multi-parallel sample for fine-tuning, the off-target problem, which has been considered a big issue, can be easily addressed (from 51.8% to 1.9%). As the fine-tuning data size is exponentially scaled up beyond 100, we observe nearly linear improvements.
@@
Due to the high efficiency and practicality, we encourage the community to consider fine-tuning with tiny readily available multi-parallel data, like NTREX (Federmann et al., 2022), as a strong baseline. Also, we call on the community to build more comprehensive multi-parallel datasets to cover real-world language types.
2 Related Works
@@
The zero-shot translation capability of MMT is revealed to be associated with multilingualism, following the hypothesis of universal representation or an interlingua.  Arivazhagan et al. (2019) view zero-shot translation as a domain adaptation problem (Ben-David et al., 2006) in MMT, and apply auxiliary losses to explicitly incentivize the model to learn and use domain- (language-) invariant representation.  Liu et al. (2021) attribute the low quality of zero-shot MT to the positional correspondence to input tokens, which hinders modeling language-agnostic representation.  Pan et al. (2021) use a contrastive loss to close the representation gap between different languages. Some other studies aim to harness the capabilities of pretrained multilingual models for zero-shot translation.  Chen et al. (2022) employ multilingual pretrained encoders to extend bilingual translation to many-to-one translation, relying on the encoder’s language-agnostic representation. Recently, some works have focused on leveraging pre-trained large language models for multilingual translation (Zhang et al., 2023; Moslem et al., 2023). Despite the inclusion of the so-called "emergent abilities" (Wei et al., 2022) triggered by zero-shot prompting, we categorize these works as following a similar line.
@@
This paper focuses on a data-centric approach for comprehensively improving zero-shot performance. We empirically show that a well-trained English-centric model can be easily boosted for overall zero-shot capability via fine-tuning with minimal multi-parallel data, even if only covering a small set of translation directions (10%). This allows us to leverage multi-parallel data which is hard to obtain in large quantities, leading to a highly efficient and practical mean of overall zero-shot translation. Note that Maillard et al. (2023) also show that incorporating small high-quality data (a few thousand samples) for training can have a big impact on low-resource translation systems. However, we argue that the reasons for the effectiveness differ: 1) our method can work with extremely minimal fine-tuning data (100 or even single samples), and 2) as shown in Section 3.5, the evident enhancements persist when using data built from the training set, meaning that the influence from domain or quality level is eliminated.
@@ 3.1 Datasets
Europarl [^2] https://www.statmt.org/europarl (Koehn, 2005) consists of 20 English-centric language pairs from the proceedings of the European Parliament, with sizes ranging from 399K to 2M. A characteristic of the Europarl dataset is that part of the samples are multi-way aligned. In this paper, we select the most resource-rich 8 languages, i.e., EN, DA, DE, ES, FI, FR, IT, and NL, to mine a fully multi-parallel dataset named Europarl-8 via aligning multi-way sentences with the same English part. As a result, we get around 1.2M fully multi-parallel data, where each sentence has 7 counterparts in other languages.
@@
To ensure a more diverse and inclusive large-scale evaluation, we follow  (Tan and Monz, 2023; Wu and Monz, 2023) and use the EC30 dataset, which is built from WMT (Bojar et al., 2017) and OPUS (Tiedemann, 2012) corpora. The EC30 dataset consists of 61 million English-centric bilingual sentences as training data, covering 30 non-English languages across a wide spectrum of resource availability, ranging from High (5M) to Medium (1M), and Low (100K) resources. Each resource group includes languages from 5 families with multiple writing systems.
@@
For all of the experiments in this paper, we evaluate translations via the Flores-101 benchmark (Goyal et al., 2022). Flores comprises 3001 sentences sourced from English Wikipedia, which covers a variety of topics and domains and is translated into 101 languages by professional translators. We use dev and devtest as the validation and test dataset, consisting of 997 and 1012 samples, respectively. All results are evaluated on three widely used metrics, i.e., ChrF++ (Popović, 2017), SacreBLEU (Post, 2018)333nrefs:1—case:mixed—eff:no—tok:13a—smooth:exp—version:2.3.1, and COMET (Rei et al., 2020), to demonstrate the consistency of improvements across a broad spectrum of evaluation metrics.
A more detailed description of datasets is provided in Appendix A.1.
@@ 3.2.1 Training Setting
For the experiments on the EC30 dataset, we use Transformer-Big with 16 attention heads, 1,024 embedding dimensions, and 4,096 feedforward dimensions. For Europarl-8, we use a smaller backbone considering less data involved, where a standard 6-layer encoder 6-layer decoder transformer model is applied with 4 attention heads, 512 embedding dimensions, and 1,024 feedforward dimensions.
For all of the English-centric training, the learning rate is 5e-4 with 4,000 warmup steps and a inverse sqrt decay schedule. All dropout rates and label smoothing are set to 0.1. In the case of EC30 and Europarl-8, the batch size is set as 8196 tokens, accumulating gradients 20 and 8 times, respectively. Also, data from different language pairs are sampled with a temperature of 5.0 and 2.0, respectively. The same temperature is applied to both BPE building and MMT training periods. We train all models with an early-stopping strategy [^4] Patience is set to {10, 20}, i.e., training stops if performance on the validation set does not improve for the last {10, 20} checkpoints, with 1,000 steps between checkpoints. and evaluate by using the best checkpoint as selected based on the loss on the development set.
@@ 3.2.2 Fine-tuning Setting
We use standard fine-tuning for all of the experiments, resetting all of the running statuses, including optimizer, lr scheduler, and data loaders. Also, the fine-tuning parameters are aligned with those in the training period, except for the experiments in Analysis (Section 4), where we set batch accumulation as 1 as super-low fine-tuning data is involved. Note that although previous works show that fine-tuning is parameter-sensitive, we keep our setup as simple as possible to highlight generalizability.
@@
Table 2: English-centric performance (ChrF) on the EC30 dataset (61M sentence pairs). EN-X and X-EN denote the average out-of- and into-English translation performance of each resource group, respectively. Results in SacreBLEU and COMET are provided in Table 13 and Table 15, respectively.
Figure 2: Zero-shot performance (ChrF) on EC30 in each scaling step, grouped by High-, Medium, and Low-resource, respectively. (a) When we randomly selected {10%, 20%, 40%, 80%} of directions to boost, the overall zero-shot performance nearly stayed unchanged. However, (b) when we fixed the 10% of directions and extended the fine-tuning samples from 100 to 800, consistent improvements can be observed for all of the resource groups.
@@ 3.3 Large-Scale Experiments on EC30
In this section, we show how far a tiny amount of multi-parallel data can boost the zero-shot capability of an already well-trained large-scale English-centric MMT. We conduct experiments on EC30, involving 30 English-centric and 870 zero-shot directions. We build fine-tuning data based on NTREX to cover all except for Occitan-related directions (Occitan is not included in NTREX), as described in Section 3.1. It is noteworthy that MMT systems typically use two language tag versions: 1) the one-tag version, i.e., putting the target language IDs on the encoder side (Johnson et al., 2017), as demonstrated by  Wu et al. (2021) that to be more effective for zero-shot capability, and 2) the two-tag version, i.e., putting source and target language IDs on the encoder and decoder, respectively, which is applied to recent large-scale MMT systems (Fan et al., 2021; Pan et al., 2021; Costa-jussà et al., 2022). To show comprehensive results, we conducted experiments in both settings.
@@
Table 1 shows the zero-shot performance across 9 resource groups, considering all combinations for source and target resource levels. Boost-All means using all of the 1997 multi-parallel samples from NTREX to construct pair-wise fine-tuning data for all directions (including English-centric ones), while Boost-100 means only using 100 randomly sampled ones, instead of 1997, to construct.
@@
We find that 1) fine-tuning with tiny data leads to very evident overall improvements for both tagging versions, with up to +9.3 and +22.7 average ChrF point gains, respectively. 2) The zero-shot capability of the two-tag baseline lags behind the one-tag baseline, in line with the finding in  (Wu et al., 2021). However, after fine-tuning with multi-parallel data, the overall performance in the two-tag setting consistently outperforms the one-tag setting for each group, yielding an average margin of +1.5 ChrF (35.5 v.s., 34.0). Consistent improvements also hold for other metrics (see Appendix A.2).
@@
In Table 2, we show the impact that fine-tuning has on English-centric directions: The trade-off effect mainly occurs in the high-resource group. However, the influence upon the medium- and low-resource groups is negligible or even positive (especially for the low-resource part), resulting in nearly unchanged overall English-centric performance. For example, by fine-tuning with 100 multi-parallel samples on the two-tag model, we achieve +21.7 ChrF zero-shot gains with negligible English-centric overall drops (-0.1 ChrF). In Table 6, we show that 854 out of 870 zero-shot directions get evident boosts (more than 10.0 ChrF).
@@
It is noteworthy that fine-tuning with just 100 samples achieves comparable improvements to using the entire NTREX dataset (+21.7 v.s., +22.7 in Table 1), although the latter’s size is 20 times larger. The pronounced diminishing marginal effect naturally leads to the following questions for the scale effects of the need for fine-tuning data:
  • What is more important, more fine-tuning data or more fine-tuning directions?
  • If we have a lot of multi-parallel data, how closely can our method approach the upper bound improvements?
@@
We try to answer these questions in the following Section 3.4 and Section 3.5. If not specified, we utilize the two-tag version model for the subsequent experiments since the facts of its higher absolute zero-shot and English-centric performance after fine-tuning compared to the one-tag version, as well as a more pronounced discrepancy compared to the corresponding baseline.

### 26 Feb 2024 v2

# How Far can 100 Samples Go? Unlocking Zero-Shot Translation with Tiny Multi-Parallel Data

@@ Abstract
Zero-shot translation aims to translate between language pairs not seen during training in Multilingual Machine Translation (MMT) and is largely considered an open problem. A common, albeit resource-consuming, solution is to add as many related translation directions as possible to the training corpus. In this paper, we show that for an English-centric model, surprisingly large zero-shot improvements can be achieved by simply fine-tuning with a very small amount of multi-parallel data. For example, on the EC30 dataset, we obtain up to +21.7 ChrF non-English overall improvements (870 directions) by using only 100 multi-parallel samples while preserving English-centric translation quality. When investigating the size effect of fine-tuning data and its transfer capabilities, we found that already a small, randomly sampled set of fine-tuning directions is sufficient to achieve comparable improvements. The resulting non-English performance is close to the complete translation upper bound. Even in a minimal setting—fine-tuning with only one single sample—the well-known off-target issue is almost completely resolved, explaining parts–but not all—of the observed improvements in translation quality. [^1] https://github.com/research-anonymous/MultiParallelFinetuning4MMT
@@ 1 Introduction
The zero-shot capability shown by Multilingual Machine Translation (MMT) Johnson et al. (2017) is of considerable significance, particularly in the context of translating between low-resource or distant language pairs. However, even for systems trained on large-scale data, the zero-shot performance is still far from sufficient (Tan and Monz, 2023), especially when scaling up the number of involved languages. Substantial efforts (Zhang et al., 2020; Pan et al., 2021; Gu and Feng, 2022; Mao et al., 2023) have been dedicated to improving the zero-shot capabilities of models trained on readily available, predominantly English-centric corpora.
@@
Figure 1: (a) English-centric training data is normally readily available but can only cover a few real-world directions, while (b) complete translation (Freitag and Firat, 2020) aims to cover all but suffers from the small data scale. (c) Mining partial non-English data as the bridge languages shows promising zero-shot improvements but is also resource-consuming when scaling up. (d) We show that substantial overall improvements can be achieved by fine-tuning an English-centric model with tiny extra multi-parallel data, which is readily available, like NTREX (Federmann et al., 2022).
@@
To fully cover translation directions, Freitag and Firat (2020) propose to mine multi-parallel (multi-way aligned) examples to extend the training set from English-centric to a complete multilingual one as shown in Figure 1-(b). Non-English translation quality in this setting indeed increases substantially. However, such a setting is far away from real-world practice when scaling up. As shown in Freitag and Firat (2020), to solely extend the training set from bilingual aligned to all languages (6-way) involved in their case, the amount of available data drops from 123M to 10K, which is insufficient.
@@
To reflect translation needs worldwide, Fan et al. (2021) build and open-source a training dataset covering 100 languages through industry-scale mining. In addition to English-centric data, supervised data for thousands of bridge language pairs is mined and included, organized based on language families. The MMT model trained on the resulting data, M2M100, exhibits clear non-English improvement in many non-English directions. This work drives a simple but also resource-consuming solution for real-world demand: mining as much training data as possible to bridge non-English language pairs, at least on the language family level.
@@
In this paper, we take a step back and again look at the readily available English-centric model. We empirically show that the corresponding zero-shot ability can be easily unlocked via fine-tuning an English-centric model with a tiny amount of multi-parallel data, which is much simpler and more efficient than the extensive bridge data mining done by earlier work. Furthermore, we investigate the size effect of fine-tuning data: 1) Surprisingly, even when fine-tuning using randomly sampled 10% of directions, the overall improvements are comparable to that of full-direction fine-tuning. 2) The improvements brought by very small fine-tuning datasets only slightly lag behind the upper bound (complete translation) while preserving English-centric capabilities, showing great practical potential. 3) Even with just one single multi-parallel sample for fine-tuning, the well-known off-target problem (Zhang et al., 2020; Yang et al., 2021; Sennrich et al., 2023), is easily addressed, reducing the off-target rate from 51.8% to 1.9%. However, not all improvements in translation quality can be solely attributed to lower off-target rates as we also see clear improvements in cases where translations are already in the correct target language.
@@
Due to the high efficiency and practicality, we encourage the community to consider fine-tuning with tiny readily available multi-parallel data, like NTREX (Federmann et al., 2022), as a strong baseline for zero-shot translation.
2 Related Work
@@
The zero-shot translation capability of MMT is associated with multilingualism, following the hypothesis of universal representation or interlingua. Arivazhagan et al. (2019) view zero-shot translation as a domain adaptation problem (Ben-David et al., 2006) in MMT, and apply auxiliary losses to explicitly incentivize the model to learn and use domain- (language-) invariant representation. Liu et al. (2021) attribute the low quality of zero-shot MT to the positional correspondence to input tokens, which hinders modeling language-agnostic representation. Pan et al. (2021) use a contrastive loss to close the representation gap between different languages. Some other approaches aim to harness the capabilities of pretrained multilingual models for zero-shot translation. Chen et al. (2022) employ multilingual pretrained encoders to extend bilingual translation to many-to-one translation, relying on the encoder’s language-agnostic representation. Recently, some work has focused on leveraging pre-trained large language models for multilingual translation (Zhang et al., 2023; Moslem et al., 2023). Despite the inclusion of the so-called “emergent abilities” (Wei et al., 2022) triggered by zero-shot prompting, we categorize these works as following a similar line.
@@
This paper focuses on a data-centric approach for comprehensively improving zero-shot performance. We empirically show that a well-trained English-centric model can be easily boosted for overall zero-shot capability via fine-tuning with minimal multi-parallel data, even if only covering a small set of translation directions (10%). This allows us to leverage multi-parallel data which is hard to obtain in large quantities, leading to a highly efficient and practical means for overall zero-shot translation. We note that Maillard et al. (2023) also show that integrating small high-quality data (6K samples) into the training corpus can have a big impact on low-resource translation systems, especially when combined with back translation (Sennrich et al., 2016). However, we argue that the reasons for the effectiveness differ: 1) as shown in Section 3.6, the substantial enhancements persist when using data built from the training set, meaning that the influence from domain or quality level is eliminated, and 2) our method can work with extremely minimal fine-tuning data (100 or even a single sample).
@@ 3 Experiments
In this paper, we propose a simple approach that leverages small amounts of multi-parallel data to fine-tune an English-centric model for a large number of directions. The fine-tuning data is constructed from small, readily available multi-parallel datasets, like NTREX (Federmann et al., 2022). We refer to such a process as “fine-tuning with multi-parallel data” or “multi-parallel fine-tuning”.
3.1 Fine-Tuning Data Construction
Given a multi-parallel dataset comprising $N$ distinct languages, each with $K$ samples, we can generate pairwise data in all $N\times(N-1)$ possible directions. Note that acquiring large quantities of multi-parallel data poses challenges due to many professional human translators being involved. However, horizontally expanding a readily available multi-parallel dataset to include one more language is straightforward. It simply requires annotating $K$ additional samples for the new language based on the current dataset, with $2\times N$ new translation directions indirectly covered.
@@ 3.2 Datasets
Europarl [^3] https://www.statmt.org/europarl (Koehn, 2005) consists of 20 English-centric language pairs from the proceedings of the European Parliament, with sizes ranging from 399K to 2M. A characteristic of Europarl is that part of the samples are multi-way aligned. In this paper, we select the most resource-rich 8 languages, i.e., EN, DA, DE, ES, FI, FR, IT, and NL, to mine a fully multi-parallel dataset named Europarl-8 via aligning multi-way sentences with the same English part. This results in about 1.2M fully multi-parallel data instances, where each sentence has 7 counterparts in other languages.
@@
To ensure a more diverse and inclusive large-scale evaluation, we follow Tan and Monz (2023); Wu and Monz (2023) and use the EC30 dataset, which is built from WMT (Bojar et al., 2017) and OPUS (Tiedemann, 2012) corpora. EC30 comprises 61 million English-centric bilingual sentences for training, encompassing 30 non-English languages with diverse resource levels (High: 5M, Medium: 1M, Low: 100K). Each resource group includes languages from 5 families with multiple writing systems.
@@
For all of the experiments in this paper, we evaluate translations via the Flores-101 benchmark (Goyal et al., 2022). Flores comprises 3001 sentences sourced from English Wikipedia, which covers a variety of topics and domains and is translated into 101 languages by professional translators. We use dev and devtest as the validation and test dataset, consisting of 997 and 1012 samples, respectively. All results are evaluated on three widely used metrics, namely, ChrF++ (Popović, 2017), SacreBLEU (Post, 2018)444nrefs:1—case:mixed—eff:no—tok:13a—smooth:exp—version:2.3.1, and COMET (Rei et al., 2020), to demonstrate the consistency of improvements across a broad spectrum of evaluation metrics. A more detailed description of the datasets is provided in Appendix A.1.
@@ 3.3.1 Training Setting
For experiments on the EC30 dataset, we use Transformer-Big with 16 attention heads, 1,024 embedding dimensions, and 4,096 feedforward dimensions. For Europarl-8, we use a smaller backbone, as the training data is smaller, where a standard 6-layer encoder, 6-layer decoder transformer model is applied with 4 attention heads, 512 embedding dimensions, and 1,024 feedforward dimensions. In total, 447M and 64M training parameters are involved for the two models. More detailed training settings are provided in Appendix A.2.
@@ 3.3.2 Fine-Tuning Setting
We use full-parameter fine-tuning and keep our setup as simple as possible to highlight generalizability. We reset all running statuses, including optimizer, lr scheduler, and data loaders. Also, the fine-tuning parameters are aligned with those in the training period, except for the experiments in Section 4, where we set batch accumulation as 1 as extremely small fine-tuning data is used.
@@
Table 2: English-centric performance (ChrF) on the EC30 dataset (60 directions, 61M sentence pairs). EN-X and X-EN denote the average out-of- and into-English translation performance of each resource group, respectively. Results in SacreBLEU and COMET are provided in Table 13 and Table 15, respectively.
Figure 2: Zero-shot performance (ChrF) on EC30 for each scaling step, grouped by High-, Medium, and Low-resource, respectively. (a) When we randomly selected {10%, 20%, 40%, 80%} of fine-tuning directions, overall zero-shot performance nearly stayed unchanged. However, (b) when we fixed 10% of directions and increased the fine-tuning samples from 100 to 800, consistent improvements can be observed for all resource groups.
@@ 3.4 Large-Scale Experiments on EC30
In this section, we show how far a tiny amount of multi-parallel data can improve the zero-shot capability of an already well-trained large-scale English-centric MMT system. We conduct experiments on EC30, involving 30 English-centric and 870 zero-shot directions. We build fine-tuning data based on NTREX to cover all except Occitan-related directions—Occitan is not included in NTREX—as described in Section 3.2. It is noteworthy that MMT systems typically use two language tag strategies: 1) the one-tag strategy, i.e., adding the target language IDs to the encoder input, which is shown by Wu et al. (2021) to be more effective for zero-shot translation, or 2) the two-tag strategy, i.e., adding source and target language IDs to the encoder and decoder input, respectively, which is often applied to recent large-scale MMT systems (Fan et al., 2021; Pan et al., 2021; Costa-jussà et al., 2022). To show comprehensive results, we conducted experiments in both settings.
@@
Table 1 shows the zero-shot performance across 9 resource groups. Boost-All means using all of the 1,997 multi-parallel samples from NTREX to construct pair-wise fine-tuning data for all directions (including English-centric ones), while Boost-100 means only using 100 randomly sampled samples, instead of 1,997, to construct the fine-tuning data.
@@
We find that 1) fine-tuning with tiny data leads to very strong overall improvements for both tagging strategies, with up to +9.3 and +22.7 average ChrF point gains, respectively. 2) The zero-shot capability of the two-tag baseline lags behind the one-tag baseline, in line with Wu et al. (2021). However, after fine-tuning with multi-parallel data, the overall performance in the two-tag setting consistently outperforms the one-tag setting for each group, yielding an average margin of +1.5 ChrF (35.5 v.s., 34.0). Consistent improvements also hold for other metrics, see Appendix A.3.
@@
In Table 2, we show the impact of fine-tuning on English-centric directions: The trade-off effect mainly occurs in the high-resource group. However, the influence on the medium- and low-resource groups is negligible or even positive, especially for the low-resource part, resulting in nearly unchanged overall English-centric performance. For instance, fine-tuning with 100 multi-parallel samples on the two-tag model yields +21.7 ChrF zero-shot gains, with negligible drops in averaged English-centric performance (-0.1 ChrF). In Table 6, we show that 854 out of 870 zero-shot directions get strong boosts (more than 10.0 ChrF).
@@
It is noteworthy that fine-tuning with just 100 samples achieves comparable improvements to using the entire NTREX dataset (+21.7 v.s. +22.7 in Table 1), even though the latter’s size is 20 times larger. This diminishing effectiveness naturally leads us to ask (i) whether more fine-tuning data or more fine-tuning directions is important and (ii) how close can our method come to the upper-bound improvements.
@@
We answer both questions in Section 3.5 and 3.6, respectively. If not specified, we employ the two-tag strategy in subsequent experiments because of its higher zero-shot and English-centric performance after fine-tuning.